# LInK: Learning Joint Representations of Design and Performance Spaces through Contrastive Learning for Mechanism Synthesis

### v1

@@ 3 Methodology
Our methodology is illustrated in Figure 1, where we begin by retrieving potential matches from a dataset of 10 million mechanisms using a processed version of the input curve. These initial matches seed our optimization algorithm to refine the designs further. This section elaborates on our dual approach. Firstly, we train a contrastive learning model that bridges the physics underlying the mechanisms with their design representations. This model is crucial for accurately identifying relevant mechanisms from the extensive dataset. Secondly, we detail our optimization algorithm, termed the Learning-accelerated Inverse Kinematics (LInK) algorithm, which builds upon the foundations laid by the contrastive retrieval model. This ’deep optimization’ strategy enhances our capability to fine-tune and improve the designs based on the initial matches identified by the retrieval process.
@@ 3.1 Contrastive Learning Framework
The first stage of our framework involves retrieving candidates from a massive dataset of linkage mechanisms for a given target curve such that these candidates are best suited to be refined for the given input. We use the LINKS dataset proposed by Heyrani Nobari et al. (2022), which includes 100 million mechanisms with up to 20 joints. Searching through such a massive dataset of this size by comparing curves and using conventional methods would be prohibitively expensive and time-consuming. Furthermore, most conventional metrics for searching in such datasets are sensitive to rotations, scale, and other spatial transformations, which makes them less useful. In our approach to facilitate this, we build a cross-modal contrastive embedding to quickly search through a massive dataset in practical time. In this section, we describe the details of how we built this model and how we build robustness to different spatial and geometric transformations.
3.1.1 Mechanism Representation & Model
@@
We employ a subset of 10 million mechanisms from the LINKS dataset, which undergoes extensive preprocessing for normalization. To achieve scale invariance, we standardize the actuator arm length to 0.05 units. We further normalize these mechanisms with respect to translation by centering the actuators at the origin. Additionally, the mechanisms are aligned such that the actuator angle is set to zero, standardizing both position and rotation. This normalization process is critical for effective training of our contrastive learning model. Furthermore, we meticulously process the dataset to eliminate any redundant joints i.e., there is one joint in the mechanism whose kinematic solution depends on the kinematics of all the other joints.
@@
In our approach, we model planar linkage mechanisms as undirected graphs $M$ as we described prior with slight variation. The only difference in the graphs we use for training is that instead of only two joint types and initial positions, the node features of the graph are defined such that The first element row indicates if a node is fixed or moving (i.e. 0 for moving joints and 1 for fixed joints), the second element indicates whether a joint is actuated and the third indicates if a given joint is the terminal joint (i.e., the joint which should produce the target curve), and the two other elements in each row are set to be the initial positions of the nodes at time zero. In training our model, we employ a sophisticated multi-layer, multi-level architecture inspired by the Graph Isomorphism Network (GIN) approach proposed by  Xu et al. (2019). Off-the-shelf graph models proved inadequate due to the intricate relationship between mechanism configurations and their kinematics, which standard message-passing techniques struggle to capture. To address this, we devised a novel architecture, which we term Graph Isomorphism Hop-Attention (GHop), inspired by mechanism solvers. This architecture enables both shallow and deep message passing within the same graph structure, optimizing computational efficiency. Our approach modifies the traditional GIN computation as follows:
@@
where $h_{v}^{(k)}$ is the output features for node/joint $v$ at layer $k$ and $\epsilon^{(k)}$ is a learnable parameter as described in the original work by Xu et al. (2019) and $\mathcal{N}(v)$ refers to the set of neighbors of node $v$. In our approach, we introduce hop-attention, where at the end of computing all layers we compute a final feature for each node through a graph attention (GAT) driven attention mechanism. To do this we take the outputs of all layers concatenate them for each node and perform a single GAT convolution on the resulting graph:
$h_{v}^{(GAT)}=\underset{n=1}{\overset{n=N}{\|}}\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{n}\mathbf{W}^{n}\underset{k=1}{\overset{k=K}{\|}}h_{v}^{(k)}\right),$ (4)
where $h_{v}^{(GAT)}$ is the output of the graph attention convolution and $N$ is the number of attention heads in the GAT and $W^{n}$ is the weights of each attention head linear transformation while $\sigma$ is a non-linearity function and $h_{v}^{(k)}$ is the output of the $k$-th layer of the GIN model. Then we use this output of the GAT as the query and the outputs of each layer for a given node as the keys and values for a multi-headed dot-product attention which is applied to obtain the final output of each layer. This means that at the end of each node/joint in the graph/mechanism, we have the following:
@@
where $h_{v}$ is the final output for node/joint $v$ and $N$ is the number of attention heads and $\left[h_{v}^{(GAT)}\right]$ is a single row matrix with $h_{v}^{(GAT)}$ as its only row and $\left[h_{v}^{(1)}h_{v}^{(2)}\cdots h_{v}^{(M)}\right]$ is a matrix whose columns are assembled by the output of each layer of the GIN model. $\mathbf{W}^{Q}_{i},\mathbf{W}^{K}_{i},\mathbf{W}^{V}_{i}$ refer to the weight of the linear transformation applied at each attention head, and $\mathbf{W}^{O}$ is the final linear transformation applied to the concatenated attention head outputs resulting the final representation for each node/joint. The intuition behind this kind of architecture is evident in the solver’s approach to computing the kinematics of the mechanisms. In the solver we start with the known joints (i.e.  0-1 hops), then propagate the solution to a joint with two known neigbours (1st hop) and repeat this process until all joints are solved (variable number of hops depending on the mechanism skeleton). This means to correctly describe the kinematics of the mechanism each joint requires a different number of hops (see figure 2) and conventional graph convolution approaches simply do not account for this as the nature of problems they encounter are not like the ones we have in our problem. Given this, it is intuitive to add a mechanism to account for the variable number of hops needed for each joint, and our GHop approach enables this through a simple yet effective attention mechanism. One thing to note is the hop attention mechanism (HAT) is independent of the graph convolution mechanism used and any graph convolution or message passing approach can be used with HAT, which should enable improvement in existing models in problems that involve similarly hop-count sensitive challenges (see figure 3 for details on the architecture).
@@
3.1.2 Path Representation & Building Robust Invariances for Retrieval
@@
  • Timing Invariance: We also need to build invariance to timing; i.e., the curves should be made up of points that are equidistant from one another. To do this, we simply take any given curve and interpolate points across the curves such that the points are equidistant.
@@
The curves in the LINKS dataset are always closed. However, the target curves that one can give to a model during testing may not always be closed. To effectively manage both closed and open target curves in our contrastive cross-modal space, we implement a robust preprocessing strategy that accommodates varying curve types. To do this, we randomly generate partial curves of the original curves that a mechanism generates, normalize them exactly like we do for the full curves, and remove timing from them (see Figure 3). We then introduce partial curves as a separate module into the overall contrastive learning space to ensure that open curves are also handled by the contrastive representation space. This is a unique contribution of this approach, as previous approaches are limited to closed curves.
@@ 3.1.3 Consolidated Methodology
In this section, we consolidate our methodologies for training contrastive learning models and searching within the dataset. Our approach maps both mechanisms and curves into a unified representation space using contrastive learning. This mapping allows any introduced curve to be accurately paired with corresponding mechanism candidates based on its spatial representation. We employ a dual-model strategy for curves alongside a specialized graph-based model for mechanisms. Specifically, the mechanism model utilizes the previously discussed GI-HAT architecture. For curve processing, we utilize two ResNet50 models (He et al., 2015): one handling full curves and another for partial curves. The training leverages two distinct contrastive losses, akin to the methodology used in the CLIP model (Radford et al., 2021):
$\mathcal{L}_{\text{LInK}}=\mathcal{L}_{\text{CLIP1}}+\gamma\mathcal{L}_{\text{CLIP2}}$ (7)
where $\mathcal{L}_{\text{CLIP1}}$ is the loss for the contrastive signal between the mechanism models output, and $\mathcal{L}_{\text{CLIP2}}$ is the contrastive loss for the partial and full curves models, and $\gamma$ is a hyperparameter for applying different weights to these losses.
@@
$\mathcal{L}_{\text{CLIP2}}$ $=\frac{1}{N}\sum_{i=1}^{N}-\log\frac{\exp(\text{sim}(g_{i},h_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(g_{i},h_{j})/\tau)}$ (9)
where $\text{sim}(a,b)$ measures the cosine similarity between feature embeddings $a$, and $b$ extracted by any of the models, and $\tau$ is the temperature parameter controlling the sharpness of the similarity scores. Importantly, the output of the mechanism model corresponds to $f_{i}$, the output of the full curve model corresponds to $g_{i}$, and the output for the partial curve model corresponds to $h_{i}$, and $N$ is the batch size. Notably, mechanisms are matched to their corresponding curves, presumed dissimilar to others in the batch. While this assumption holds given our diverse dataset, no specific measures have been implemented to handle potential similarities between different mechanisms’ curves. This setup is visualized in figure 3, detailing the interaction between these components in our training framework.
@@ 3.2 Path Synthesis Framework
    Batch Optimization: In the second stage, we perform a gradient-based optimization based on the BFGS BROYDEN (1970); Fletcher (1970); Goldfarb (1970); Shanno (1970) method upon all of the mechanisms retrieved in the first stage. To accelerate the process, we implement a batch BFGS optimization on GPU which performs the BFGS optimization on multiple mechanisms simultaneously. We describe this in more detail in this section. In this stage we only perform 10 steps of optimization as optimizing the larger initial pool of candidates can be slow.
@@
    Final Refinement: After the batch optimization of the candidates, we then re-evaluate the mechanisms and pick the top 10% of mechanisms amongst the candidates to move to the next stage. In the final stage of optimization, we further refine the smaller candidate pool using the BFGS optimization for 150 steps. at which point the final design is output to the user.
@@
3.2.1 GPU-Accelerated Batch Simulation of Linkage Mechanisms & Differentiation
@@
We write a solver to simulate mechanisms, similar to the one proposed in Heyrani Nobari et al. (2022) for dataset generation. This solver handles the kinematics of mechanisms geometrically. While it is not suitable for the rare instances of mechanisms with complex kinematic loops—a limitation discussed further in the prior work Heyrani Nobari et al. (2022)—it suits our needs. The advantage of this solver is twofold. The entire process is differentiable with respect to the initial positions of the joints in a given mechanism and the process can be vectorized for a batch of simulations which makes this solver the perfect candidate for a differentiable GPU accelerated framework like ours. Here we will briefly discuss the solver and how we vectorized the solver for batches of simulations.
@@
To vectorize and parallelize this process we have to perform two things:
  • Pre-Sorting Mechanisms: To facilitate vectorization and batch processing, we first establish a consistent solution path for all mechanisms in the dataset. By pre-sorting the joints of each mechanism in a solution-ready order, we eliminate the need to identify the solution path in real time during simulations. This sorting, conveniently, has already been handled by Heyrani Nobari et al. (2022) method, as proposed by the creators.
@@
With the mechanisms appropriately sorted and resized, the next step is to implement a vectorized solver that can efficiently handle batches of mechanisms. This solver operates on the principle of solving for any given joint based on the positions of two known adjacent joints at each timestep using the followingequation Bächer et al. (2015):
$\mathbf{X}_{i}^{T}=\mathbf{X}_{j}^{T}+||X_{i}^{T}-X_{j}^{T}||\times\mathbf{R}(\phi)\frac{X_{k}^{T}-X_{j}^{T}}{||X_{k}^{T}-X_{j}^{T}||}$ (10)
@@
$\begin{split}\phi=&\text{ Sign}\left[\left(X_{j,y}^{0}-X_{i,y}^{0}\right)\times\left(X_{j,x}^{0}-X_{k,x}^{0}\right)-\left(X_{j,y}^{0}-X_{k,y}^{0}\right)\times\left(X_{j,x}^{0}-X_{i,x}^{0}\right)\right]\times\\ &\cos^{-1}{\frac{||X_{j}^{T}-X_{k}^{T}||^{2}+||X_{i}^{T}-X_{j}^{T}||^{2}-||X_{i}^{T}-X_{k}^{T}||^{2}}{2\times||X_{j}^{T}-X_{k}^{T}||^{2}\times||X_{i}^{T}-X_{j}^{T}||}}\end{split}$ (11)
where $\mathbf{X}_{i}^{0}$ is the initial positions of joint $i$. As can be seen, these equations are easily vectorizable for multiple mechanisms (essentially by turning the $X$ vectors into tensors with an additional dimension with the same size as the batch size) and multiple time steps (all $T$s that need to be solved for). Furthermore, all of the above-mentioned equations are differentiable (except the sign function which will not be differentiable at zero), which allows us to use backpropagation to compute the gradients with respect to the initial positions $X^{0}$ and run the simulation much faster in a vectorized fashion for a batch of mechanisms and for all timesteps simultaneously. We implement such a GPU-accelerated solver using Pytorch and use the automatic differentiation features of Pytorch to obtain the gradients of the solution with respect to the initial positions of the joints.
@@
$d_{CD}\left(S_{1},S_{2}\right)=\frac{1}{|S_{1}|}\sum_{x\in S_{1}}\min_{y\in S_{2}}\|x-y\|_{2}+\frac{1}{|S_{2}|}\sum_{y\in S_{2}}\min_{x\in S_{1}}\|x-y\|_{2}$ (12)
@@
where $N$ is the number of points sampled in both the target curve and mechanism traced curve and $X^{\text{coupler}}_{i}$ is the $i$-th point on the curve traced by the mechanisms and $X^{\text{target}}_{i}$ is the $i$-th point on the target curve. This while $o_{1}(i)$ is the $i-th$ point in the optimal ordering to match the curves and $O_{1}$ is the set of all possible clockwise and counter-clockwise orderings of the curve traced by a mechanism. The constant $2\pi$ is kept for consistency across different works to allow for meaningful comparison of results to prior works that use this metric. Finally, in our gradient-based optimization, we have the following objective function:
$d=\gamma_{1}\times d_{OD}+\gamma_{2}\times d_{CD},$ (14)

### v2

@@ 3 Methodology
Our methodology is illustrated in Figure 1, where we begin by retrieving potential matches from a dataset of 10 million mechanisms using a processed version of the input curve. These initial matches seed our optimization algorithm to refine the designs further. This section elaborates on our dual approach. Firstly, we train a contrastive learning model that bridges the physics underlying the mechanisms with their design representations. This model is crucial for accurately identifying relevant mechanisms from the extensive dataset. Secondly, we detail our optimization algorithm, termed the Learning-accelerated Inverse Kinematics (LInK) algorithm, which builds upon the foundations laid by the contrastive retrieval model. This ‘deep optimization’ strategy enhances our capability to fine-tune and improve the designs based on the initial matches identified by the retrieval process.
@@ 3.1 Contrastive Learning Framework
The first stage of our framework involves retrieving candidates from a massive dataset of linkage mechanisms for a given target curve such that these candidates are best suited to be refined for the given input. We use the LINKS dataset proposed by Heyrani Nobari et al. (2022), which includes 100 million mechanisms with up to 20 joints. Searching through such a massive dataset of this size by comparing curves and using conventional methods would be prohibitively expensive and time-consuming. Furthermore, most conventional metrics for searching in such datasets are sensitive to rotations, scale, and other spatial transformations, which makes them less useful. To improve this in our approach, we build a cross-modal contrastive embedding to quickly search through a massive dataset in practical time. In this section, we describe the details of how we built this model and how we build robustness to different spatial and geometric transformations.
3.1.1 Mechanism Representation and Model
@@
We employ a subset of 10 million mechanisms from the LINKS dataset, which undergoes extensive preprocessing for normalization. To achieve scale invariance, we standardize the actuator arm length to 0.05 units. We further normalize these mechanisms with respect to translation by centering the actuators at the origin. Additionally, the mechanisms are aligned such that the actuator angle is set to zero, standardizing both position and rotation. This normalization process is critical for the effective training of our contrastive learning model. Furthermore, we meticulously process the dataset to eliminate any redundant joints, i.e., a joint in the mechanism whose kinematic solution depends on the kinematics of all the other joints.
@@
In our approach, we model planar linkage mechanisms as undirected graphs $M$ as we described above with slight variation. The only difference between the graphs we use for training and the prior described graphs is that instead of only two joint types and initial positions, the node features of the graph are defined such that the first element row indicates if a node is fixed or moving (0 for moving joints and 1 for fixed joints), the second element indicates whether a joint is actuated and the third indicates if a given joint is the terminal joint (i.e., the joint which should produce the target curve). The two other elements in each row are set to be the initial positions of the nodes at time zero. In training our model, we employ a sophisticated multi-layer, multi-level architecture inspired by the Graph Isomorphism Network (GIN) approach proposed by  Xu et al. (2019). Off-the-shelf graph models proved inadequate due to the intricate relationship between mechanism configurations and their kinematics, which standard message-passing techniques struggle to capture. To address this, we devised a novel architecture which we term Graph Isomorphism Hop-Attention (GHop), inspired by mechanism solvers. Unlike traditional GNNs where message passing occurs uniformly across all layers, GHop allows for adaptive processing of nodes. This is particularly relevant in mechanism synthesis, where nodes are only processed until they are solved, as illustrated in Figure 5. GHop enables the model to adapt the receptive field for each node separately through an attention mechanism, allowing it to incorporate this feature of linkage mechanisms and their solutions in an unsupervised way. This architecture enables both shallow and deep message passing within the same graph structure, optimizing computational efficiency and improving the representation of linkages compared to baseline GNN models. Our approach modifies the traditional GIN computation as follows:
@@
where $h_{v}^{(k)}$ is the output features for node/joint $v$ at layer $k$, $\epsilon^{(k)}$ is a learnable parameter as described in the original work by Xu et al. (2019) and $\mathcal{N}(v)$ refers to the set of neighbors of node $v$. In our approach, we introduce hop-attention, where at the end of computing all layers we compute a final feature for each node through a graph attention (GAT) driven attention mechanism. To do this, we take the outputs of all layers, concatenate them for each node, and perform a single GAT convolution on the resulting graph:
$h_{v}^{(GAT)}=\underset{n=1}{\overset{n=N}{\|}}\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}^{n}\mathbf{W}^{n}\underset{k=1}{\overset{k=K}{\|}}h_{v}^{(k)}\right),$ (4)
where $h_{v}^{(GAT)}$ is the output of the graph attention convolution, and $N$ is the number of attention heads in the GAT. $W^{n}$ is the weights of each attention head linear transformation, while $\sigma$ is a nonlinearity function, and $h_{v}^{(k)}$ is the output of the $k$-th layer of the GIN model. We then use the output of the GAT as the query. We additionally use the outputs of each layer for a given node as the keys and values for a multi-headed dot-product attention. The attention is applied to obtain the final output of each layer. This means that at the end of each node/joint in the graph/mechanism, we have the following:
@@
where $h_{v}$ is the final output for node/joint $v$ and $N$ is the number of attention heads. $\left[h_{v}^{(GAT)}\right]$ is a single row matrix with $h_{v}^{(GAT)}$ as its only row, and $\left[h_{v}^{(1)}h_{v}^{(2)}\cdots h_{v}^{(M)}\right]$ is a matrix whose columns are assembled by the output of each layer of the GIN model. $\mathbf{W}^{Q}_{i},\mathbf{W}^{K}_{i},\mathbf{W}^{V}_{i}$ refer to the weight of the linear transformation applied at each attention head, and $\mathbf{W}^{O}$ is the final linear transformation applied to the concatenated attention head outputs resulting the final representation for each node/joint. The intuition behind this kind of architecture is evident in the solver’s approach to computing the kinematics of the mechanisms. In the solver we start with the known joints (i.e.,  0-1 hops), then propagate the solution to a joint with two known neighbors (1st hop) and repeat this process until all joints are solved (variable number of hops depending on the mechanism skeleton). This means to correctly describe the kinematics of the mechanism each joint requires a different number of hops (see Figure 5). Conventional graph convolution approaches simply do not account for this, as the nature of the problems they encounter are not like the ones we have in our problem. Given this limitation, it is intuitive to add a mechanism to account for the variable number of hops needed for each joint. Our GHop approach enables this through a simple yet effective attention mechanism. One thing to note is the hop attention mechanism (HAT) is independent of the graph convolution mechanism used and any graph convolution or message passing approach can be used with HAT, which should enable improvement in existing models to solve problems that involve similarly hop-count sensitive challenges (see Figure 4 for details on the architecture).
@@
3.1.2 Path Representation and Building Robust Invariances for Retrieval
@@
  • Timing Invariance: We also need to build invariance to timing; i.e., the curves should be made up of points that are equidistant from one another. To do this, we simply take any given curve and interpolate points across the curves such that the points are equidistant. Being invariant to timing and making paths equidistant is important for path synthesis since in this kind of problem we do not care about the speed and timing of the motion rather just the shape (see Appendix C for more details).
@@
The curves in the LINKS dataset are always closed. However, the target curves that one can give to a model during testing may not always be closed. To effectively manage both closed and open target curves in our contrastive cross-modal space, we implement a robust preprocessing strategy that accommodates varying curve types. To do this, we randomly generate partial curves of the original curves that a mechanism generates, normalize them exactly like we do for the full curves, and remove timing from them (see Figure 4). To gather these samples we randomly cut out a portions of the full closed curves (between 10% to 100% of the curve) and process them as described before. We then introduce partial curves as a separate modality into the overall contrastive learning space to ensure that open curves are also handled by the contrastive representation space. This is a unique contribution of this approach, as previous approaches are limited to closed curves. The ability to handle partial curves is particularly challenging and important. Unlike closed curves, partial curves lack a clear correspondence with the full mechanism-generated curves. This makes traditional matching techniques, such as Procrustes analysis, ineffective as they rely on known correspondences. Moreover, the scale and translation of partial curves relative to the full curves are unknown, further complicating the matching process. By incorporating partial curves into our contrastive learning framework, we enable the model to effectively capture and match these challenging cases, significantly expanding the applicability of our approach to a broader range of real-world scenarios.
@@ 3.1.3 Consolidated Methodology
In this section, we consolidate our methodology for training contrastive learning models and searching within the dataset. Our approach maps both mechanisms and curves into a unified representation space using contrastive learning. This mapping allows any introduced curve to be accurately paired with corresponding mechanism candidates based on its spatial representation. We employ a dual-model strategy for curves alongside a specialized graph-based model for mechanisms. Specifically, the mechanism model utilizes the previously discussed GHop architecture. For curve processing, we utilize two ResNet50 models (He et al., 2015): one handling full curves and another for partial curves. The training leverages two distinct contrastive losses, akin to the methodology used in the CLIP model (Radford et al., 2021):
$\mathcal{L}_{\text{LInK}}=\mathcal{L}_{\text{CLIP1}}+\gamma\mathcal{L}_{\text{CLIP2}},$ (7)
where $\mathcal{L}_{\text{CLIP1}}$ is the loss for the contrastive signal between the mechanism model’s output, and $\mathcal{L}_{\text{CLIP2}}$ is the contrastive loss for the partial and full curves models. $\gamma$ is a hyperparameter for applying different weights to these losses.
@@
$\mathcal{L}_{\text{CLIP2}}$ $=\frac{1}{N}\sum_{i=1}^{N}-\log\frac{\exp(\text{sim}(g_{i},h_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(g_{i},h_{j})/\tau)},$ (9)
where $\text{sim}(a,b)$ measures the cosine similarity between feature embeddings $a$ and $b$ extracted by any of the models, and $\tau$ is the temperature parameter controlling the sharpness of the similarity scores. Importantly, the output of the mechanism model corresponds to $f_{i}$, the output of the full curve model corresponds to $g_{i}$, and the output for the partial curve model corresponds to $h_{i}$, and $N$ is the batch size. Notably, mechanisms are matched to their corresponding curves, and presumed dissimilar to others in the batch. While this assumption holds given our diverse dataset, no specific measures have been implemented to handle potential similarities between different mechanisms’ curves. Handling partial curves as a separate modality is crucial due to their inherently noisier signal. Partial curves often contain common repeated patterns, such as arcs, which can lead to ambiguous associations if directly matched with mechanisms. By treating them separately and matching them with full curves, we leverage more detailed geometric information like curvature, allowing for cleaner associations. This approach helps maintain the quality of the learned representations and prevents the introduction of noise into the mechanism-curve associations. This setup is visualized in Figure 4, detailing the interaction between these components in our training framework.
@@ 3.2 Path Synthesis Framework
    Batch Optimization: In the second stage, we perform a gradient-based optimization building off of the BFGS BROYDEN (1970); Fletcher (1970); Goldfarb (1970); Shanno (1970) method, which is applied to all of the mechanisms retrieved in the first stage. To accelerate the process, we implement a batch BFGS optimization on GPU which performs the BFGS optimization on multiple mechanisms simultaneously. We describe this in more detail in this section. In this stage we only perform 10 steps of optimization as optimizing the larger initial pool of candidates can be slow.
@@
    Final Refinement: After the batch optimization of the candidates, we then re-evaluate the mechanisms and pick the top 10% amongst the candidates to move to the next stage. In the final stage of optimization, we further refine the smaller candidate pool using the BFGS optimization for 150 steps, at which point the final design is output to the user.
@@
3.2.1 GPU-Accelerated Batch Simulation of Linkage Mechanisms and Differentiation
Figure 5: This figure illustrates the path the solver takes to solve the kinematics of a given mechanism, showing how the skeleton of a mechanism (combinatorial design variables) is involved in the mechanism solution. Initially, the solver starts with the known joints (i.e., fixed and actuated joints highlighted red), and step by step the solver solves joints with two solved neighbors (Eqn. 10).In this example, joint 3 (the last joint to be solved) is solved in three steps. The numbered joints indicate the order of solution.
@@ 3.2.1 GPU-Accelerated Batch Simulation of Linkage Mechanisms and Differentiation
We write a solver to simulate mechanisms, similar to the one proposed in Heyrani Nobari et al. (2022) for dataset generation. This solver handles the kinematics of mechanisms geometrically. While it is not suitable for the rare instances of mechanisms with complex kinematic loops—a limitation discussed further in the prior work Heyrani Nobari et al. (2022)—it suits our needs. The advantage of this solver is twofold. First, the entire process is differentiable with respect to the initial positions of the joints in a given mechanism and second, the process can be vectorized for a batch of simulations which makes this solver the perfect candidate for a differentiable GPU accelerated framework like ours. Here we will briefly discuss the solver and how we vectorized the solver for batches of simulations.
@@
To vectorize and parallelize this process we have to perform two steps:
  • Pre-Sorting Mechanisms: To facilitate vectorization and batch processing, we first establish a consistent solution path for all mechanisms in the dataset. By presorting the joints of each mechanism in a solution-ready order, we eliminate the need to identify the solution path in real time during simulations. This sorting, conveniently, has already been handled by Heyrani Nobari et al. (2022) method.
@@
With the mechanisms appropriately sorted and resized, the next step is to implement the vectorized solver that can efficiently handle batches of mechanisms. This solver operates on the principle of solving for any given joint based on the positions of two known adjacent joints at each timestep using the following equation Bächer et al. (2015):
$\mathbf{X}_{i}^{T}=\mathbf{X}_{j}^{T}+||X_{i}^{T}-X_{j}^{T}||\times\mathbf{R}(\phi)\frac{X_{k}^{T}-X_{j}^{T}}{||X_{k}^{T}-X_{j}^{T}||},$ (10)
@@
$\begin{split}\phi=&\text{ Sign}\left[\left(X_{j,y}^{0}-X_{i,y}^{0}\right)\times\left(X_{j,x}^{0}-X_{k,x}^{0}\right)-\left(X_{j,y}^{0}-X_{k,y}^{0}\right)\times\left(X_{j,x}^{0}-X_{i,x}^{0}\right)\right]\times\\ &\cos^{-1}{\frac{||X_{j}^{T}-X_{k}^{T}||^{2}+||X_{i}^{T}-X_{j}^{T}||^{2}-||X_{i}^{T}-X_{k}^{T}||^{2}}{2\times||X_{j}^{T}-X_{k}^{T}||^{2}\times||X_{i}^{T}-X_{j}^{T}||}},\end{split}$ (11)
where $\mathbf{X}_{i}^{0}$ is the initial position of joint $i$. As can be seen, these equations are easily vectorizable for multiple mechanisms (essentially by turning the $X$ vectors into tensors with an additional dimension of the same size as the batch size) and multiple time steps. Furthermore, all of the above-mentioned equations are differentiable (except the sign function which will not be differentiable at zero). This allows us to use backpropagation to compute the gradients with respect to the initial positions $X^{0}$ and run the simulation much faster in a vectorized fashion for a batch of mechanisms and for all timesteps simultaneously. We implement such a GPU-accelerated solver using Pytorch and use the automatic differentiation features of Pytorch to obtain the gradients of the solution with respect to the initial positions of the joints.
@@
$d_{CD}\left(S_{1},S_{2}\right)=\frac{1}{|S_{1}|}\sum_{x\in S_{1}}\min_{y\in S_{2}}\|x-y\|_{2}+\frac{1}{|S_{2}|}\sum_{y\in S_{2}}\min_{x\in S_{1}}\|x-y\|_{2},$ (12)
@@
where $N$ is the number of points sampled in both the target curve and mechanism traced curve. $X^{\text{coupler}}_{i}$ is the $i$-th point on the curve traced by the mechanisms, and $X^{\text{target}}_{i}$ is the $i$-th point on the target curve. This while $o_{1}(i)$ is the $i$-th point in the optimal ordering to match the curves and $O_{1}$ is the set of all possible clockwise and counter-clockwise orderings of the curve traced by a mechanism. The constant $2\pi$ is kept for consistency across different works to allow for meaningful comparison of results to prior works that use this metric. Finally, in our gradient-based optimization, we have the following objective function:
$d=\gamma_{1}\times d_{OD}+\gamma_{2}\times d_{CD}.$ (14)